OpenStudy (anonymous):

Can Someone Help On 7&8? ILL GIVE YOU A MEDAL<3

OpenStudy (anonymous):

Whats the question ?

OpenStudy (anonymous):

Ohhh , thats hard . You should be able to answer it , you the smart one lol

OpenStudy (anonymous):

No i HATE MATH REMEMber!:p

OpenStudy (anonymous):

LOL I do too , remember ?

OpenStudy (anonymous):

i kno who can help me!:) maybe @gorv

OpenStudy (anonymous):

Alright then :) lol sorry i couldnt help

OpenStudy (anonymous):

no problammo

OpenStudy (gorv):

okkk

OpenStudy (gorv):

initially height of ball will be =0

OpenStudy (anonymous):

ok

OpenStudy (gorv):

height is given that means distance of the ball from ground

OpenStudy (anonymous):

h(t)

OpenStudy (gorv):

yeah velocity=distance/time

OpenStudy (gorv):

or v=d(x)/dt....where x is distance and function of time

OpenStudy (gorv):

diffeentiate h(t) to get speed

OpenStudy (anonymous):

idk i give up /; i hate this sh..t>;(

OpenStudy (gorv):

okk ...how abt dis vellocity = dx/dt if x is function of time here h is function of time so differntiate it to get velocity v=dh/dt=-16*2*t+32=-32t+32 at highest point speed of ball will become zero 0=-32t+32 32t=32 t=1 sec

OpenStudy (the_fizicx99):

Lol I'm pretty sure you're suppose to use the quadratic formula :)

OpenStudy (gorv):

nope

OpenStudy (anonymous):

so its a? for 7

OpenStudy (gorv):

ok now put t=1 in h(t) to get max height h(1)=-16*(1)^2+32*1+6=-16+32+6=22 meter

OpenStudy (gorv):

yeah moni its A

OpenStudy (anonymous):

and 8

OpenStudy (gorv):

similarly do for 8 differntiate h(t) 0=-32t+184 t=184/32=5.75 sec

OpenStudy (anonymous):

its d or c

OpenStudy (anonymous):

*Grabs Popcorn* Reads thru comments.

OpenStudy (gorv):

put t in h(t)=-16*5.75^2+184*5.75+20 =549

OpenStudy (anonymous):

lol @JavierCerda

OpenStudy (gorv):

C

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

thanks